If the Perimeter = 6x+6 and x=width, find the binomial representing the length.
Answer by checkley75(3666)

6X+6=2X+2L
6X-2X+6=2L
2L=4X+6
L=2X+3 ANSWER FOR THE LENGTH.
PROOF
2X+2(2X+3)=6X+6
2X+4X+6=6X+6
6X+6=6X+6